▾▴ Toggle Texts
▾▴ Toggle Images
◱ Full Size Pics
⊻ Bottom of Page
- (inserted: 10.03.2006 22:00)
The Peten Itza Scientific Drilling Project has finished the lastThe Peten Itza Scientific Drilling Project has finished the last hole and site of the project. Without enough time to move to an additional site, the team finished drilling today with 1327 meters of core in the reefer. Demobilization will begin shortly. We would like to thank everyone involved in the project for their support. Each of you has helped make this project successful.
- (inserted: 09.03.2006 22:00)
The first and second holes at site Pi-6 were finishedThe first and second holes at site Pi-6 were finished today with an average depth of 68 meters below lake floor and 97% recovery. The team will drill one last hole at the site for the conclusion of the PISDP.
- (inserted: 08.03.2006 22:00)
New core liner arrived on the platform for the dayNew core liner arrived on the platform for the day shift. The shift returned with over 50 meters of sediment.
- (inserted: 07.03.2006 22:00)
Operations are still on hold while we wait for coreOperations are still on hold while we wait for core liners to arrive.
- (inserted: 06.03.2006 22:00)
Today packing of the MSCL equipment and the laboratory equipmentToday packing of the MSCL equipment and the laboratory equipment was finalized. Drilling is still paused pending the arrival of new core liners.
- (inserted: 05.03.2006 22:00)
The DOSECC team continues with platform maintainence.The DOSECC team continues with platform maintainence.
- (inserted: 04.03.2006 22:00)
The platform was towed to Pi-6.The platform was towed to Pi-6. Because of a lack of core liner, drilling will be postponed briefly, and rig maintenance will be performed. The project has begun to wind down with the departure of Flavio this afternoon. The team plans to continue drilling with a reduced crew through the 13th of March.
- (inserted: 03.03.2006 22:00)
Today the platform was towed to Pi-9 and drilling began.Today the platform was towed to Pi-9 and drilling began. A group also collected modern day surface sediment samples along a depth transect. The team encountered gravely deposits at the site. Today one of our drilling team members, Ryan left for medical reasons. We would like to thank Ryan for all of his hard work during the project.
- (inserted: 02.03.2006 22:00)
On Mark's birthday, we finished site Pi7 which is ourOn Mark's birthday, we finished site Pi7 which is our longest record (140m) !
- (inserted: 01.03.2006 22:00)
The team celebrated Dave's birthday today after finishing the secondThe team celebrated Dave's birthday today after finishing the second hole at Pi-7.
- (inserted: 28.02.2006 22:00)
Wednesday, the PISDP recovered the 1000th meter of sediment fromWednesday, the PISDP recovered the 1000th meter of sediment from Peten Itza.
- (inserted: 27.02.2006 22:00)
The final depth of hole A was 133 meters belowThe final depth of hole A was 133 meters below lake floor. This depth was a very exciting surprise to everyone. Mardi Gras was celebrated in el Remate. Coring at Pi-7 B began with the evening shift change.
- (inserted: 25.02.2006 22:00)
The Scientific Advisory Group of the ICDP visited the PISDPThe Scientific Advisory Group of the ICDP visited the PISDP at La Casa de Don David on Sunday. Progress reports were given and the group toured the field lab. Coring at Pi-7 started on the Sunday/Monday night shift. The group will visit the platform on Monday to see the Glad 800 in action.
- (inserted: 24.02.2006 22:00)
Logging at the site through the drill string was successful.Logging at the site through the drill string was successful. The ICDP logging crew was also partially able to log in the open hole. The team will prepare the barge for towing overnight. The R/V Kerry Kelts be towed to Pi-7 on Sunday.
- (inserted: 23.02.2006 22:00)
The first hole at site Pi-4 was stopped at aThe first hole at site Pi-4 was stopped at a depth of 65 meters below lake floor. Within the cores there are several mixed layers of gravel and gypsum which are thought to be caused by a combination of turbidites and/or slumping at this hole. The second hole was started Friday afternoon. The borehole logging team from the GFZ plans to log the second hole Saturday morning.
- (inserted: 21.02.2006 22:00)
After a day of towing and rig maintainence the teamAfter a day of towing and rig maintainence the team celebrated Doug's birthday with a great cake made by the staff at La Casa de Don David. Drilling at the deepest site (Pi-4) is expected to begin Thursday.
- (inserted: 20.02.2006 22:00)
This night we scored a hat-trick with the completion ofThis night we scored a hat-trick with the completion of the third site, PI-3. Wednesday morning the barge will be towed to Pi-4, our deepest site.
- (inserted: 18.02.2006 22:00)
As the night shift neared completion, we finished drilling theAs the night shift neared completion, we finished drilling the first hole at site Pi-3, reaching bedrock at a depth of about 100 m. The day shift began drilling of the second hole at the site, recovering 33 m of core with almost 100% recovery.
- (inserted: 17.02.2006 22:00)
Drilling at the third site (Pi-3) began Friday night.Drilling at the third site (Pi-3) began Friday night. The night and day crews have again settled into a working groove. Recovery from this first hole has been excellent (98%.) Based on the past rate of progress at the site the team hopes to reach bedrock within the next 24 hours. The seismic profile suggests bedrock will be roughly 100 meters below lake floor.
- (inserted: 16.02.2006 22:00)
The platform was moved successfully to Pi-3 where drilling isThe platform was moved successfully to Pi-3 where drilling is expected to begin during the night shift Friday. Some team members took the day to catch up on email, and others surveyed the area for raised lake deposits.
- (inserted: 15.02.2006 22:00)
The third and final hole at site Pi-1 was completedThe third and final hole at site Pi-1 was completed Thursday afternoon. The DOSECC team will prepare the platform for towing to Pi-3 Friday morning.
- (inserted: 14.02.2006 22:00)
Several students from the Universidad del Valle joined the teamSeveral students from the Universidad del Valle joined the team at el Remate. The students have been very excited to participate in the project and plan to stay for several days before leaving to continue work on other lakes in the region. The downhole logging team was successful with logging efforts through the casing and directly within the borehole.
- (inserted: 13.02.2006 22:00)
Progress at the second site (Pi-1) has been swift.Progress at the second site (Pi-1) has been swift. With the drilling techniques optimized, recovery at the second hole (B) was excellent (94%.) The logging team from the ICDP is expected to log the B hole during the night shift. Cores have been streaming into the reefer.
- (inserted: 12.02.2006 22:00)
The night shift recovered a solid piece of bedrock limestoneThe night shift recovered a solid piece of bedrock limestone at 4:00AM Monday morning while using the "Alien" tool. The depth below lake floor was approximately 95 Meters. The night and day drilling teams have both been doing an exceptional job . Visiting scientists have provided a steady supply of expatriated snacks and coffee which have been greatly appreciated.
- (inserted: 11.02.2006 22:00)
Cores arrived at el Remate with the night shift team,Cores arrived at el Remate with the night shift team, and more arrived as the day shift returned. The operations are going smoothly and the teams have enjoyed the suprisingly cooler (55F/13C) temperatures.
- (inserted: 10.02.2006 22:00)
With great help from Franz Sperisen of the Camino RealWith great help from Franz Sperisen of the Camino Real Tikal the R/V Kerry Kelts was towed to the second site on Saturday. The DOSECC day crew put down the casing and prepared the platform for the night shift to start drilling soon after the shift change. Cores are expected to arrive from the shore-base on Sunday morning.
- (inserted: 09.02.2006 22:00)
The Ortiz sits at el Remate, awaiting a new motor.The Ortiz sits at el Remate, awaiting a new motor. The loss of our towing boat has delayed operations for a day and had threatened to delay the movement of the platform for several more days. With all other options explored Dave and Flavio visited the Camino Real Tikal and spoke with Franz Sperisen about the possibility of using their large vessel for towing. With great generosity and personal and professional risk Franz happily agreed to allow the vessel to be used for towing. The towing will begin Saturday morning at 6:00 AM. As this news was announced at the shore-base the team morale rocketed. Franz was really our saving grace.
- (inserted: 08.02.2006 22:00)
The PISDP team has begun preparations to move the platformThe PISDP team has begun preparations to move the platform to the second drillng site. A film crew from Swiss TV has started filming the team for a science magazine (MTW).
- (inserted: 07.02.2006 22:00)
Despite some minor obstacles the first site is finished. TheDespite some minor obstacles the first site is finished. The spliced density data suggest that the team has collected a complete composite core to the depth of 65 meters. Of particular interest were unexpected layers of elemental sulfur. We look forward to the next cores as we move the platform to site Pi-1.
- (inserted: 06.02.2006 22:00)
With the second hole completed the ICDP logging team beganWith the second hole completed the ICDP logging team began measurements. Initial measurements through the drill string provided some interesting natural gamma data. Unfortunately the second measurements taken after tripping (or pulling) the drill string were not as successful. We suspect that this was due to an unstable hole.
- (inserted: 05.02.2006 22:00)
The team experienced a setback as a seal blew onThe team experienced a setback as a seal blew on our hydraulic hoist. After several members of the DOSECC team worked a +24 hour shift, drilling resumed Monday. The correlation of the first two holes is being carried out at the shore-base using the MSCL data and Splicer.
- (inserted: 04.02.2006 22:00)
Drilling operations during the night shift reached 41 meters atDrilling operations during the night shift reached 41 meters at the second hole. The drilling teams and the lab teams are finally getting used to their shifts. The logging team from the ICDP/GFZ are on site and preparing for their deployment tomorrow.
- (inserted: 03.02.2006 22:00)
The day-shift team hit bedrock early this afternoon.The day-shift team hit bedrock early this afternoon. Coring began on the second hole.
- (inserted: 30.01.2006 22:00)
With the platform in the water a beautiful sunset isWith the platform in the water a beautiful sunset is even prettier.
- (inserted: 26.01.2006 22:00)
Anticipitation rises as drilling platform nears completion.Anticipitation rises as drilling platform nears completion.
- (inserted: 22.01.2006 22:00)
Today, at six o`clock in the evening - what aToday, at six o`clock in the evening - what a great moment! Our material has arrived at the border of Guatemala/Belize. The picture shows a huge truck with material for the GLAD-800 drilling platform.
- (inserted: 18.01.2006 22:00)
The PISDP drilling and science teams are all at theThe PISDP drilling and science teams are all at the Casa de Don David. We have set up a field laboratory. We are still awaiting the arrival of the drilling platform which we hope will occur during the next few days.
- (inserted: 15.12.2005 22:00)
The Peten project is moving forward with drilling equipment inThe Peten project is moving forward with drilling equipment in Guatemala. Two graduate students have just finished the GFZ DIS training. Flavio Anselmetti, Dave Hodell and the rest of the scientific team are getting excited for the drilling to start. Grüezi mitänand :-)